OpenStudy (anonymous):

how do i solve this system equations by using the row reduction method? {2x+y=2 3x+2y=7

OpenStudy (goformit100):



OpenStudy (anonymous):

are yu sure?

OpenStudy (goformit100):

yes

OpenStudy (goformit100):

:)

OpenStudy (anonymous):

okay thanks :)

OpenStudy (richyw):

\[\left[\begin{matrix}2 & 1 &2 \\ 3 & 2 & 7\end{matrix}\right]\]Multiply row 1 by 3/2 \[\left[\begin{matrix}3 & \frac{3}{2} &3 \\ 3 & 2 & 7\end{matrix}\right]\]Subtract row one from row 2 \[\left[\begin{matrix}3 & \frac{3}{2} &3 \\ 0 & \frac{1}{2} & 4\end{matrix}\right]\]this gives \(\frac{y}{2}=4 \) so \(y=8\) \[3x+\frac{3}{2}y=3\]\[3x+\frac{24}{2}=3\]\[3x=-9\]\[x=-3\]